HUBBLE SPACE TELESCOPE DAILY REPORT #1011 PERIOD COVERED: 0800 EST Daily Status Report as of 313/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed HSP/POL 3250 (Gravitational Lenses Part 1 - Cycle 2 High) The High Speed Photometer was used to make ultraviolet photometric observations of quasar QSO0957+56 which is an expected gravitational lenes system. It is hoped to find a characteristic signature in variability in the time scale of the order of weeks. The observations were monitored, were possible. Most of the observations were received in real-time and displayed. The data were reported to be normal. The onboard target acquisition was successful. The remainder of the data were received from the science tape recorder, and they were reported to look okay. 1.2 Completed FGS 5068 (Linking the HIPPARCOS to the Extragalactic Reference Frame part 6 of 6) The Fine Guidance Sensor #3 was used to make two sets of astrometric observations of object NEWEGOD-54 and reference e object 220-REF. This object lies near the apparent position of extragalactic objects. This object will also be observed with the HIPPARCOS satellite. The intent is to tie the HIPPARCOS system reference frame to the extragalactic reference frame. The observations were executed as scheduled. The observation occurred during LOS and could not be monitored in real-time. The observation of the reference object 220-REF was monitored in real-time and was reported to look okay. 1.3 Completed FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star system HD181391 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The continuous scans were monitored and no problems were seen in the telemetry. 1.4 Completed GHRS/2 5041 (GHRS Sensitivity Monitor - Cycle 3) The High Resolution Spectrograph was used to make ultraviolet observations of standard star BD28D4211 to monitor the sensitivity of the HRS, which is expected to change slowly with time. The onboard target acquisition was successful. Both sets of observations were received and displayed in the OSS. The target was well centered. The astrometric observations were reported to look okay. 1.5 Completed GHRS/2 5024 (GHRS Telescope Breathing Monitor - Cycle 3) The High Resolution Spectrograph was used to make observations of the standard star HD184499 to look for variations in the flux through the Small Science Aperture (SSA) and shifts in position during the course of several orbits, which may be caused by "breathing" in the secondary mirror. The observations were monitored in the OSS, and the data quality was assessed. No anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 6 There was a Guide Star Acquisition failure at 313/0445Z during a ZOE. A search limit was exceeded for FGS #1. An ARU/PRT was performed and the subsequent reacquisition was successful. Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Significant HSTARs: None 2.4 Operations Note: Simulation JIS #6 (EVA 4 and 5, plus Deploy) was initiated yesterday morning, Monday, Nov. 8, 1993 at 8 AM and is scheduled to last for 59 hours. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 4510 (The Evolution of Boron: Implementation for Galactic Evolution) 3.2 Complete GHRS/2 2536 (Deuterium in the Local Interstellar Gas) 3.3 Complete WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.4 Complete Two Sets fo WFC 5076 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 2) 3.5 Complete Two Sets of PC 5072 (WF/PC Cycle 3 Calibration: PC Flats Wide. 1) 3.6 Complete Two Sets of PC 5073 (WF/PC cycle 3 Calibration: PC flats Wide. 2) 3.7 Complete GHRS/2 3472 (The Properties of Single Interstellar Clouds - Cycle 2) 3.8 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-